OpenStudy (anonymous):

What is the point slope form of (0, –2), (3, 2)?

OpenStudy (anonymous):

\[m=\frac{ y2-y1 }{ x2-x1 }\] does it help?

OpenStudy (anonymous):

 Is this what it would look like?

OpenStudy (anonymous):

Yes, that would give you the slope. Then you pick one of the points and that slope and substitute the x-value of the point you choose for x1 and the y-value of the point you choose for y1 into: y-y1 = m(x-x1) This then gives y - 2 = 4/3 (x-3) or y -(-2) = 4/3 (x-0) which simplifies to y + 2 = 4/3 x

OpenStudy (anonymous):